# Ergodicity detection algorithms: Scaling of ergodicity in random symbolic dynamics

### v1

@@ Abstract
ergodic hypothesis; randomness; symbolic dynamics;ensemble theory
pacs:
MSC:, 37A60,82B03, 82B41, 82M31, ACM:, G.3, G.2.3, I.6.1
I Introduction
@@
The origins of the ergodic hypothesis appear to be statements put forward by Boltzmann and Maxwell, that multiple states evolving with constant energy should return to their initial conditions [1]. This simple-sounding statement as the origin of Poincaré recurrance [2] formed the basis of the connection between statistical mechanics and its thermodynamic justification, formalised by Gibbs ensemble theory [3]. However, the interpretation that a system should visit all possible states as a mapping from constant energy phases or from nearby points (quasi-ergodic) [4] to its time evolution attracted the interest of mathematical scientists [5], i.e., von Neumann and Birkhoff via their theorems [6, 7, 8], in some sense diverging from statistical physics to form a mathematical field called ergodic theory for low dimensional systems as an extension of strong form of law of large numbers [9, 10, 11]. In this context, the term ergodicity, a probe that measures if a given dynamical regime attains ensemble averages being the same as time averages or if this equivalence is broken, is studied as a key aspect of the dynamics of a physical system [12].
@@
Measuring ergodicity is practiced and needed in many different applied quantitative fields, such as, simple liquids [13], the Ising-Lenz model with external fields and Hopfield networks [14, 15], earthquake fault dynamics [16], atmospheric forecasting [17], economics [18], financial portfolios [19], machine learning of optimal policies [20], neuromorphic computing [21], random matrices [22, 23], deep learning architecture search [24, 25], random binary sequences [26], artificial spin-ice experiments [27], relativistic geodesics [28], markov chains [29, 30, 31] and network congestion [32]. In biomedicine, sensitivity analysis in differential equations for disease trajectories would also require an attention to ergodicity [33].
@@
In section II, we introduce the basic definition of ergodicity and distinctions we use for different kinds of ergodicity via sufficiency of sparse-visits, section III details of conceptual components of detection algorithms in the discrete setting are defined, section IV presents numerical experiments with random symbolic dynamics demonstrating ergodicity detection framework in the mentioned two different ergodicity schemes and their scaling properties. Finally, in section V, we conclude the study.
II Gibbs-Boltzmann ergodicity
@@
A mathematical ensemble, as introduced by Gibbs differentiate itself from probabilistic notion of sample space. This was even noted in statistical inference literature [34], where ensemble entails a more generic way of building set of possibilities. In the context of statistical mechanics, this is the set of microstates, i.e., mechanical degrees of freedoms. It might be obvious that von Neumann-Birkhoff ergodicity may not be attainable for thermodynamic systems due to astronomical number of states to visit [35]. However, Gibbs-Boltzmann ergodicity states visiting most likely portion of the energy manifold in the microcanonical ensemble, this brings the concept of sufficiency of sparse visits, that makes ergodic hypothesis plausiable. Moreover, von Neumann-Birkhoff ergodicity would still hold for low-dimensional systems [36].
@@
Gibbs-Boltzmann ergodicity manifests at very large sparsities orders close to Avogadro’s number for macroscopic systems. This made the core of the concept of sufficiency of sparse visits, whereby very small number of visits are sufficient, compare to number of states that a system actually can attain, this is due to energy manifolds values can be constructed with very small subsets of states a system can attain [1, 13, 37]. For very low dimensional systems, the sparsity level of 1.0 corresponds to von Neumann-Birkhoff ergodicity [36]. The concept of sufficiency of sparse visits reflects the thermodynamics justification of ergodic hypothesis Boltzmann and Gibbs established with contributions from Maxwell initially [1, 3]. Coarse-graining procedure is rather distinct in representing the degrees of freedom for the system before it evolves over time, on the other hand sufficiency of sparse visits occurs over-time and it’s the property of dynamics that representation reduction occurs naturally. If a system reaches to ergodic regime in between two time points $t_{0},t_{n}]$ at the acceptable $\epsilon$, then system said to be ergodicergodic. In this sense, a process is not ergodic rather a system attains the ergodic regime at a given time interval.
Figure 1: Discrete representation of an Ensemble set with alphabet size 3 and event size 2. Symbols are visualised with black, white and gray, columns are being ensemble members.
@@ III Ergodicity detection
We define a discrete dynamical system represented by the tuple for ergodicity, $\langle\mathscr{S},\mathscr{E},\mathscr{U},\mathscr{O}\rangle$, where the set of states $\mathscr{S}\in\mathbb{N}^{n}$, states can be restricted portion of $\mathscr{N}$, ensemble $\mathscr{E}$ with members $e_{j}(\mathscr{S})$, set of dynamical rules $\mathscr{U}(t)=f_{i}(e_{j})$ with $f_{i}$ being functions of ensemble members appear as recurrence equations in evolving the system from the initial states and an observable $\mathscr{O}$ as a function of ensemble members $e_{j}$. If given dynamical rules have a stochastic component then this will be random discrete dynamical system.
@@
alphabet event C $C^{se}$ $\alpha$ $\alpha^{se}$
@@
Table 1: Scaling of approach to ergodicity for diffeernt alphabet and event sizes, Fitting the power law over $\Omega$, approach to ergodicity, $Ct^{\alpha}$, with standard error values identified with nested bootstrapping.
@@
alphabet event sparsity C $C^{se}$ $\alpha$ $\alpha^{se}$
@@
alphabet event sparsity C $C^{se}$ $\alpha$ $\alpha^{se}$
@@
IV Random symbolic dynamics
Working on a discrete setting for ergodicity detection naturally bring the discrete dynamical systems and consideration of symbolic dynamics [41]. In essence, we operate on and analyse the sequences generate by an alphabet $\mathscr{A}$, which forms the states $\mathscr{S}$ of our dynamical system. We can build ensembles over symbols using different shift maps [42]. The simplest example in this context is probably a Bernoulli process, states $\mathscr{S}=\{0,1\}$. An ensemble of sequence of length 3 for these states can be represented as black and white cells. Then the ensemble looks like in Figure 2, columns being ensemble members, the set $\{000,100,010,110,001,101,011,111\}$. This can be generated via cartesian products on the alphabet repeatedly by the sequence length. Using NumPy [43], we utilize state-of-the-art random generator called PCG [44].
Similarly, we show the ensemble of $\mathscr{S}=\{0,1,2\}$ with sequence of length of 2, colour mapped white, black and gray in Figure 1.
@@ IV Random symbolic dynamics
We compute approach to ergodicity with no sparsity, corresponding to the von Neumann-Birkhoff ergodicity, for different alphabet sizes and system lengths: 65, 58, 55, 52, 48, 43, 33, 28, 25, 22. Shown in Figure 3. The rate of convergence has a smooth ordering based on alphabet size and sequence lengths. Measuring strength of the power law has similar ordering as reported in Table 1. Doubling the alphabet length approximately doubles the strenght but we see a constant scaling exponent over all lengths, about 0.5. This constant power-law scaling indicates random nature.
@@
V Conclusions
In this work we provide a clear distinction between von Neumann-Birkhoff and Gibbs-Boltzmann ergodicities reflecting the conditions where ergodic hypothesis can be attained. Along these lines, measuring and detecting ergodic regime algorithmically is formalized for the development of family of ergodicity detection algorithms. This formalisation and understanding necessary components would help practitioners in diverse fields. We have demonstrated application of the frameworks on the random symbolic dynamics, using diverse systems and identify scaling of ergodicity in these systems.
Acknowledgments: Y. Süzen for her kind support in securing research residency and encouragements on making results and tools open to research community.

### v2

@@ Abstract
keywords:
ergodic hypothesis, randomness, symbolic dynamics, ensembles
1 Introduction
@@
The origins of the ergodic hypothesis appear to be statements put forward by Boltzmann and Maxwell, that multiple states evolving with constant energy should return to their initial conditions [1]. This simple-sounding statement as the origin of Poincaré recurrence [2] formed the basis of the connection between statistical mechanics and its thermodynamic justification, formalized by Gibbs ensemble theory [3]. However, the interpretation that a system should visit all possible states as a mapping from constant energy phases or from nearby points (quasi-ergodic) [4] to its time evolution attracted the interest of mathematical scientists [5], i.e., von Neumann and Birkhoff via their theorems [6, 7, 8], in some sense diverging from statistical physics to form a mathematical field called ergodic theory for low dimensional systems as an extension of strong form of law of large numbers [9, 10, 11]. In this context, the term ergodicity, a probe that measures if a given dynamical regime attains ensemble averages being the same as time averages or if this equivalence is broken, is studied as a key aspect of the dynamics of a physical system [12].
@@
Measuring ergodicity is practiced and needed in many different applied quantitative fields, such as, simple liquids [13], the Ising-Lenz model with external fields and Hopfield networks [14, 15], earthquake fault dynamics [16], atmospheric forecasting [17], economics [18], cognitive psychology [19, 20], financial portfolios [21], machine learning of optimal policies [22], neuromorphic computing [23], random matrices [24, 25], thermal quantum dynamics [26, 27, 28], deep learning architecture search [29, 30], random binary sequences [31], artificial spin-ice experiments [32], crystal phases [33], relativistic geodesics [34], quantum computing [35], Markov chains [36, 37, 38] and network congestion [39]. In biomedicine, sensitivity analysis in differential equations for disease trajectories would also require an attention to ergodicity [40].
@@
In the first section 2, we introduce the basic definition of ergodicity and distinctions we use for different kinds of ergodicity via sufficiency of sparse-visits, the next section 3 itemized the details of conceptual components of detection algorithms in the discrete setting are defined, section 4 on experiments presents numerical experiments with random symbolic dynamics demonstrating ergodicity detection framework in the mentioned two different ergodicity schemes and their scaling properties. Finally, in the conclusion section 5, we provide final statements of the study.
2 Gibbs-Boltzmann ergodicity
@@
A mathematical ensemble, as introduced by Gibbs differentiate itself from probabilistic notion of sample space. This was even noted in statistical inference literature [41], where ensemble entails a more generic way of building set of possibilities. In the context of statistical mechanics, this is the set of micro-states, i.e., mechanical degrees of freedoms. It might be obvious that von Neumann-Birkhoff ergodicity may not be attainable for thermodynamic systems due to astronomical number of states to visit [42]. However, Gibbs-Boltzmann ergodicity states visiting most likely portion of the energy manifold in the microcanonical ensemble, this brings the concept of sufficiency of sparse visits, that makes ergodic hypothesis plausible. Moreover, von Neumann-Birkhoff ergodicity would still hold for low-dimensional systems [43].
@@
Gibbs-Boltzmann ergodicity manifests at very large sparsities orders close to Avogadro’s number for macroscopic systems. This made the core of the concept of sufficiency of sparse visits, whereby very small number of visits are sufficient, compare to number of states that a system actually can attain, this is due to energy manifolds values can be constructed with very small subsets of states a system can attain [1, 13, 44, 45]. For very low dimensional systems, the sparsity level of 1.0 corresponds to von Neumann-Birkhoff ergodicity [43]. The concept of sufficiency of sparse visits reflects the thermodynamics justification of ergodic hypothesis Boltzmann and Gibbs established with contributions from Maxwell initially [1, 3]. Coarse-graining procedure is rather distinct in representing the degrees of freedom for the system before it evolves over time, on the other hand sufficiency of sparse visits occurs over-time and it’s the property of dynamics that representation reduction occurs naturally. If a system reaches to ergodic regime in between two time points $t_{0},t_{n}]$ at the acceptable $\epsilon$, then system said to be ergodicergodic. In this sense, a process is not ergodic rather a system attains the ergodic regime at a given time interval.
Figure 1: Discrete representation of an Ensemble set with alphabet size 3 and event size 2. Symbols are visualized with black, white and gray, columns are being ensemble members.
@@ 3 Ergodicity detection
We define a discrete dynamical system represented by the tuple for ergodicity, $\langle\mathscr{S},\mathscr{E},\mathscr{U},\mathscr{O}\rangle$, where the set of states $\mathscr{S}\in\mathbb{N}^{n}$, states can be restricted portion of $\mathscr{N}$, ensemble $\mathscr{E}$ with members $e_{j}(\mathscr{S})$, set of dynamical rules $\mathscr{U}(t)=f_{i}(e_{j})$ with $f_{i}$ being functions of ensemble members appear as recurrence equations in evolving the system from the initial states [48] (dynamical law is a mapping from current states to the new points) and an observable $\mathscr{O}$ as a function of ensemble members $e_{j}$. If given dynamical rules have a stochastic component then this will be random discrete dynamical system.
@@
A size C $C^{se}$ $\alpha$ $\alpha^{se}$
@@
Table 1: Scaling of approach to ergodicity for different alphabet and event sizes, Fitting the power law over $\Omega$, approach to ergodicity, $Ct^{\alpha}$, with standard error values identified with nested bootstrapping.
@@
A size sparsity C $C^{se}$ $\alpha$ $\alpha^{se}$
@@
A size sparsity C $C^{se}$ $\alpha$ $\alpha^{se}$
@@
4 Random symbolic dynamics
Working on a discrete setting for ergodicity detection naturally bring the discrete dynamical systems and consideration of symbolic dynamics [50]. In essence, we operate on and analyze the sequences generate by an alphabet $\mathscr{A}$, which forms the states $\mathscr{S}$ of our dynamical system. We can build ensembles over symbols using different shift maps [51]. The simplest example in this context is probably a Bernoulli process, states $\mathscr{S}=\{0,1\}$. An ensemble of sequence of length 3 for these states can be represented as black and white cells. Then the ensemble looks like in Figure 2, columns being ensemble members, the set $\{000,100,010,110,001,101,011,111\}$. This can be generated via Cartesian products on the alphabet repeatedly by the sequence length. Using NumPy [52], we utilize state-of-the-art random generator called PCG [53], implementing ergodicity detection algorithm for random symbolic dynamics [54].
Similarly, we show the ensemble of $\mathscr{S}=\{0,1,2\}$ with sequence of length of 2, color mapped white, black and gray in Figure 1.
@@ 4 Random symbolic dynamics
We compute approach to ergodicity with no sparsity, corresponding to the von Neumann-Birkhoff ergodicity, for different alphabet sizes and system lengths: 65, 58, 55, 52, 48, 43, 33, 28, 25, 22. Shown in Figure 3. The rate of convergence has a smooth ordering based on alphabet size and sequence lengths. Measuring strength of the power law has similar ordering as reported in Table 1. Doubling the alphabet length approximately doubles the strength but we see a constant scaling exponent over all lengths, about 0.5. This constant power-law scaling indicates random nature.
@@
5 Conclusions
In this work we provide a clear distinction between von Neumann-Birkhoff and Gibbs-Boltzmann ergodicities reflecting the conditions where ergodic hypothesis can be attained. Along these lines, measuring and detecting ergodic regime algorithmically is formalized for the development of family of ergodicity detection algorithms. This formalization and understanding necessary components would help practitioners in diverse fields. We have demonstrated application of the frameworks on the random symbolic dynamics, using diverse systems and identify scaling of ergodicity in these systems.
Acknowledgments
Y. Süzen for her kind support and encouragements.
Declarations
  • The author declare that he has no conflict of interest.
  • There is no data release with the work.
  • Code is released under a public repository [54].